OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 5) and (0 -4)?

TheSmartOne (thesmartone):

Slope formula is \[\frac{ y _{2} -y _{1}}{ x _{2}-x _{1}}\]

TheSmartOne (thesmartone):

So (5-(-4))/(2-0) = 9/2 is your slope.

TheSmartOne (thesmartone):

(5-(-4)=5+4=9

OpenStudy (anonymous):

Thanks ;)

TheSmartOne (thesmartone):

You are welcome.